OpenStudy (anonymous):

Calc 1: Find the derivative of y = e^x / (3 * x^(1/5))

OpenStudy (anonymous):

For this problem, you will need to make use of the quotient rule and the chain rule:\[ \mathbf{Quotient\ Rule:}\ \left(\frac f g \right)'=\frac{gf'-fg'}{g^2}\\ \mathbf{Chain\ Rule:}\ (f\circ g)'=(f'\circ g)g' \]

OpenStudy (anonymous):

First step is to use the quotient rule. Set something as f and something as g, find f' and g', then plug them all into the rule.

OpenStudy (anonymous):